### v1

# JWST spectroscopy of a blue binary cold classical Kuiper belt object

Abstract
We present observations of two binary systems within the cold classical region of the Kuiper belt—2001 XR254 and 2016 BP81—obtained with the JWST Near-Infrared Spectrograph. The measured reflectance spectrum of 2001 XR254 is characteristic of the red cold classicals, with strong features due to carbon dioxide, carbon monoxide, and methanol ices. In contrast, 2016 BP81 is a blue binary, with a water-ice-rich surface composition. The two components of the 2016 BP81 binary display identical spectral profiles, consistent with coeval formation from gravitational collapse. Through qualitative and quantitative comparisons of water-ice-rich Kuiper belt objects observed with JWST, we identify a small subclass, including 2016 BP81, that appears to differ in systematic ways from the rest of the population. The relatively deep carbon dioxide ice absorption bands and enhanced signatures of aliphatic organics suggest that objects within this subclass may have originated in a distinct formation environment from the other water-ice-rich Kuiper belt objects. The implications of our findings are discussed within the context of recent models of Kuiper belt formation and evolution.
1 Introduction
The vast population of icy bodies in the Kuiper belt traces the initial compositional landscape and subsequent dynamical processes in the outer protoplanetary disk and holds the key to understanding many fundamental aspects of early solar system history. As the detailed orbital architecture of the Kuiper belt has come into focus in the past two decades, a surprisingly complex dynamical picture has emerged (e.g., B. Gladman et al., 2008; J. Kavelaars et al., 2008; M. E. Brown, 2012; J. J. Kavelaars et al., 2020), which has in turn motivated new theories of solar system evolution. Recent models attribute the broad eccentricity and inclination distribution of Kuiper belt objects (KBOs) to outward migration of Neptune that occurred shortly after the end of planet formation (A. Morbidelli & D. Nesvorný, 2020; B. Gladman & K. Volk, 2021). Numerical simulations have successfully reproduced most dynamical features of the present-day Kuiper belt by invoking a gradual expansion of Neptune’s orbit, which led to the disruption of the primordial disk of icy planetesimals that was situated beyond 30 au (e.g., H. F. Levison et al., 2008; D. Nesvorný & A. Morbidelli, 2012; D. Nesvorný & D. Vokrouhlický, 2016).
Contemporaneously, a concerted effort has been levied toward characterizing the surface properties of KBOs. Ground- and space-based multiband photometry and low-resolution spectroscopy uncovered a stark color bimodality at both visible and near-infrared wavelengths among objects fainter than an absolute magnitude of mag, i.e., smaller than 300–400 km (e.g., W. C. Fraser & M. E. Brown, 2012; W. C. Fraser et al., 2023; N. Peixinho et al., 2012, 2015; S. C. Tegler et al., 2016; I. Wong & M. E. Brown, 2017; M. E. Schwamb et al., 2019). This bimodality separates KBOs into two subpopulations that we refer to here as blue and red.111Other works in the literature apply different binary color designations, e.g., gray/red, red/very-red. The color bimodality suggests that a sharp gradient in surface composition must have existed somewhere within the primordial transneptunian disk, possibly caused by differential sublimation and irradiation chemistry that occurred on two sides of a volatile ice retention line (E. L. Schaller & M. E. Brown, 2007; M. E. Brown et al., 2011; I. Wong & M. E. Brown, 2016).
The advent of JWST has provided the means to place the attested diversity of KBO colors within a more concrete compositional framework. Near-infrared observations of dozens of KBOs in the range mag have revealed three distinct compositional classes (M. N. De Prá et al., 2025; J. Licandro et al., 2025; N. Pinilla-Alonso et al., 2025): (1) objects with deep water-ice (H2O) bands and blue visible spectral slopes that also show minor absorptions due to carbon dioxide (CO2) ice; (2) CO2-rich objects, with colors consistent with the red visible color subpopulation, that display an exceptionally deep 4.27-µm CO2 primary band, evidence of carbon monoxide (CO) ice, and strong organic features produced from the irradiation of aliphatic hydrocarbons (e.g., methane, ethane); (3) a separate subclass of red-colored KBOs with a distinct near-infrared continuum shape that additionally show characteristic features of methanol (CH3OH).222These three compositional classes are designated by the respective monikers ‘bowl’, ‘double-dip’, and ‘cliff’ in N. Pinilla-Alonso et al. (2025). Extending the aforementioned volatile ice retention hypothesis to the three compositional classes implies the existence of two compositional gradients within the primordial Kuiper belt region. Based on the observed compositional inventories, the volatile-depleted H2O-rich KBOs are expected to have formed in the innermost region of the disk, with the CO2-rich and CH3OH-rich classes accreting at progressively larger heliocentric distances (N. Pinilla-Alonso et al., 2025).
The so-called cold classicals—KBOs on relatively low-inclination, low-eccentricity orbits that lie between the 3:2 and 2:1 mean motion resonances with Neptune at 42–47 au—challenge our current understanding of Kuiper belt formation and evolution. This population is notable for containing a high fraction (30%) of near-equal-mass binaries, including many with wide separations (K. Noll et al., 2020). Dynamical studies have demonstrated that a majority of these binary systems would have been lost had they been scattered into their present-day orbits during the outward migration of Neptune (A. H. Parker & J. J. Kavelaars, 2010; D. Nesvorný & D. Vokrouhlický, 2019). Instead, cold classicals may have formed in situ and therefore comprise a unique representative collection of planetesimals from the outermost reaches of the protoplanetary disk (D. Nesvorný et al., 2022).
Within the aforementioned compositional gradient model, one would expect in situ formation of cold classicals to exclusively produce red objects. While red objects indeed dominate the cold classical population, observations within the last decade have revealed a significant admixture of bluer objects in this region; moreover, these bluer objects are predominantly binaries (W. C. Fraser et al., 2017, 2021). These blue binaries are unlikely to have formed within the inner portion of the primordial transneptunian planetesimal disk, given the fragility of widely-separated binaries to significant dynamical excitation. Instead, it has been proposed that all blue binaries formed as binary pairs immediately inward of the present-day cold classical region, between 38 and 42 au; subsequent slow migration of Neptune’s 2:1 mean motion resonance could have gently pushed these objects out into their current orbits, consistent with the somewhat inflated inclination distribution of blue binaries when compared to the redder cold classicals (W. C. Fraser et al., 2017, 2021).
However, recent attempts to simulate this evolutionary pathway have reached an impasse (e.g., D. Nesvorný et al., 2020, 2022). While a color transition immediately inward of the current cold classical region readily produces the blue binaries, it also predicts a significantly larger proportion of bluer objects throughout the inner transneptunian planetesimal disk, particularly below 30 au, where the vast majority of the primordial mass is expected to be concentrated. As a result of Neptune’s migration, a large number of blue singletons would have been scattered outward and implanted into the cold classical population, inconsistent with the observed preponderance of binaries. A drastic reformulation of the initial conditions in the transneptunian region and the specific trajectory of Neptune’s migration would be needed to resolve this quandary. An alternative explanation posits that all cold classicals (including the blue binaries) formed in situ, with the color distinction being a consequence of the timing of formation relative to the dispersion of the gas disk (D. Nesvorný et al., 2022).
Target | (au) | (∘) | , (km) | (km) | (%/100 nm) | (UT) | (au) | (au) | (∘) | (mag) | (s) | Comments | |
2001 XR254 | 42.9 | 0.035 | 1.2 | 172, 140 | 9310 | 2024 Nov 21 05:13:26 | 43.79 | 43.45 | 1.23 | 22.3 | 2976 | ||
2016 BP81 | 43.7 | 0.074 | 4.2 | 188, 170 | 11300 | 2025 Jan 18 15:52:09 | 43.29 | 43.27 | 1.31 | 22.6 | 3793 | pointing offset | |
2003 HG57 | 43.8 | 0.038 | 2.1 | 156, 156 | 13200 | 2025 Feb 15 15:23:07 | 42.45 | 42.31 | 1.33 | 22.8 | 5019 | target missed |
Note. : orbital semimajor axis; : orbital eccentricity; : orbital inclinationl , : diameters of the primary and secondary components; : mutual separation distance; : visible spectral slope; : observation start time; : heliocentric distance at time of observation; : distance from JWST at time of observation; : phase angle of observation; : visible apparent magnitude at time of observation; : total exposure time across all four dithers. The orbital parameters and viewing geometry are computed by JPL Horizons. The visible spectral slope measurements are tabulated in W. C. Fraser et al. (2021). The sizes and mutual separation of the binary components are taken from D. Nesvorný et al. (2022).
It is evident that blue binaries are a particularly diagnostic population for validating and refining our fundamental understanding of the Kuiper belt. Obtaining a detailed picture of their surface chemistry and comparing it to the compositional classes previously identified via JWST spectroscopy has the potential to resolve the ongoing conundrum concerning the provenance of blue binaries within the cold classical region. To this end, we obtained observations of several cold classical binaries as part of Cycle 3 JWST Program #5940. Section 2 describes the target selection, observing methods, and data reduction and presents the measured reflectance spectra. Comparative characterization of the spectra and discussion of their implications are provided in Section 3.
2 Observations and Data Analysis
2.1 Observation Design
The observations were carried out using the integral field unit (IFU) on the Near-Infrared Spectrograph (NIRSpec; P. Jakobsen et al. 2022; T. Böker et al. 2023). A set of four dithered exposures was collected for each target using the PRISM grating, which covers the 0.6–5.3 µmwavelength range at an average spectral resolution of . To reduce detector readnoise, all exposures were read out with the NRSIRS2RAPID method (S. H. Moseley et al., 2010; B. J. Rauscher et al., 2012). Due to the large uncertainties in the binary orbit solutions for our targets, no timing constraints were stipulated to schedule the observations to coincide with maximum component separation.
2.2 Target Selection
When selecting targets for observation as part of our Cycle 3 JWST program, we consulted the list of cold classicals with visible spectral slope measurements published in W. C. Fraser et al. (2021). Restricting our search to confirmed binaries and adopting the spectral slope value of nm as the threshold dividing the blue and red color subpopulations, we chose the three brightest blue binaries according to their minimum apparent magnitude during the Cycle 3 observing windows. The selected targets were 2001 XR254, 2003 HG57, and 2016 BP81, which have measured spectral slopes of , , and , respectively, in units of %/100 nm. Table 1 provides additional information about the targets, including component diameters, mutual separations, and the details of our observations.
As part of the proposal preparation process, the ephemerides of each target were ingested into the Astronomer’s Proposal Tool (APT) via a query of the JPL Horizons service. An important caveat of this functionality is that the ephemerides ingestion is a one-time action and not a dynamic link. As such, changes in the ephemerides from subsequent updates to the JPL Horizons orbital solutions are not automatically transferred to the APT. Our program was initially designed during the Cycle 2 call for proposals, and we neglected to reingest the latest ephemerides prior to the scheduled observations. This oversight proved to be critical for our observations of 2003 HG57, whose predicted position changed drastically in the time since the initial ingest. As a result, the target did not fall within the field of view (FOV) of the NIRSpec IFU. The ephemerides for 2016 BP81 also evolved somewhat, but not enough to cause a failed observation, albeit the location of the target within the FOV was offset from the nominal pointing position.
2.3 Data Reduction
Data processing and spectral extraction were handled by the jwstspec pipeline333https://github.com/ianyuwong/jwstspec—a dedicated tool for customized end-to-end reduction of JWST spectroscopic datasets that was developed primarily for observations of small bodies in the Solar System. This pipeline has been employed in numerous previous analyses of NIRSpec observations (e.g., J. P. Emery et al., 2024; W. M. Grundy et al., 2024; I. Wong et al., 2024), and a full description of the overall pipeline workflow can be found in those publications. Beginning with the uncalibrated data, jwstspec passed through the first two stages of Version 1.17.1 of the official JWST pipeline (H. Bushouse et al., 2025), with necessary reference files drawn from context jwst_1322.pmap of the JWST Calibration Reference Data System (CRDS). All pipeline parameters were left at their default values, except for the nsclean step in the second stage, which was manually activated to correct for residual readnoise on the detector. The result of the pipeline processing was bias-corrected, flat-fielded, spatially-rectified, wavelength-calibrated, and flux-calibrated IFU data cubes for each dithered exposure.
To compute the centroid position of the target within the FOV, the IFU data cubes were collapsed along the wavelength axis to produce a median image of the FOV. Figure 1 shows the respective median images from the first dithered exposures of 2001 XR254 and 2016 BP81. The binary nature of the targets is clearly evident. In the case of 2016 BP81, the observations took place fortuitously near maximum mutual separation, and the two components are individually resolved at the pixel scale of the NIRSpec IFU, allowing us to extract the spectrum of each component individually with negligible contamination. The slightly offset pointing for 2016 BP81 is also apparent—a consequence of the deprecated ephemerides. The jwstspec pipeline uses an iterative two-dimensional Gaussian fit to determine the centroid position from the median image. For 2016 BP81, each component was masked when centering the spectral extraction aperture of the other. In the case of 2001 XR254, the point-spread functions (PSFs) of the two components are blended, and we did not attempt to separate the two components when computing the centroid. Nevertheless, even with a single-Gaussian fit, the measured centroid is reasonable, lying intermediate between the two components, as shown in Figure 1.
The jwstspec pipeline offers several different spectral extraction techniques. For the results presented in this paper, we utilized the empirical PSF fitting method, which is optimized for faint point-source targets and yields substantial improvements in precision over standard circular aperture extraction (e.g., I. Wong et al., 2024; N. Pinilla-Alonso et al., 2025). In short, a moving median window of a specified width is applied along the wavelength axis to generate background-subtracted template PSFs, which are then normalized and fit to the central wavelength slices within the windows to obtain the best-fit multiplicative scaling factors for the templates. The fluxes are then measured from the scaled template PSFs within a user-defined circular aperture. See the referenced papers for full details on this method. For our observations, we set the background region as all pixels outside of a box centered on the measured centroid. The moving median window width was set to 51 (i.e., from to along the wavelength axis), though adjusting this value by up to did not substantively change the resultant spectra. Uncertainties on the fluxes were estimated by weighting the error arrays with the peak-normalized PSF template model and adding in quadrature the uncertainties of all pixels within the extraction aperture.
For 2001 XR254, we selected a 4-pixel-radius circular aperture for extracting the spectra, which envelops both components. For both the primary and secondary components of the 2016 BP81 system, we used a smaller 2-pixel-radius aperture to avoid contamination. The location of the extraction apertures are shown in Figure 1.
The spectra extracted from the set of four dithered exposures were passed through a 21-pixel-wide moving median filter to remove outliers. The combined irradiance spectrum of each target was obtained by a simple mean of the four individual dithered spectra. Due to low instrument throughput and poor data quality at the extreme ends of the wavelength range, we trimmed the spectra to 0.7–5.1 um.
2.4 Accounting for Sensitivity Variations
We used NIRSpec observations of the G-type solar analog star SNAP-2 when deriving the reflectance spectra of our targets. By dividing the irradiance spectrum of an object by that of the star, processed in an analogous manner and extracted from an aperture with the same equivalent encircled flux, we self-consistently corrected for the wavelength-dependent flux losses outside of the extraction apertures, while also removing any common-mode instrumental systematics that are present in the data. In the case of the blended binary 2001 XR254, our choice of the 4-pixel-radius aperture is roughly equivalent to two 3-pixel-radius apertures centered on the two components in terms of encircled flux, and we therefore used a 3-pixel-radius aperture when extracting the stellar spectrum. For 2016 BP81, the stellar spectrum was extracted with the same 2-pixel-radius aperture as the target spectrum.
The offset of 2016 BP81 relative to the nominal pointing region in the center of the FOV is almost , which is larger than the range subtended by the 4-point dither pattern. This raises the possibility of systematic sensitivity variations across the FOV that could bias the shape of the resultant reflectance spectrum when incorporating stellar observations in which the source was centered in a different region of the FOV. To investigate this issue, we considered two observations of SNAP-2, one that employed a 20-point dither pattern (Program #1128, Observation 2) and another that used the 4-point nod pattern (Program #4498, Observation 19).444For more on NIRSpec IFU dither and nod patterns, see https://jwst-docs.stsci.edu/jwst-near-infrared-spectrograph/nirspec-operations/nirspec-dithers-and-nods/nirspec-ifu-dither-and-nod-patterns. In the former case, the spread in pointing positions is similar to the 4-point dither pattern that was selected for our KBO targets, while in the latter, the inter-nod separations are much larger, with the star placed at the vertices of a box centered on the middle of the FOV.
Figure 2 shows the irradiance spectra of SNAP-2 from the two observations, both extracted using a 3-pixel-radius aperture. Also plotted is the CALSPEC model spectrum of SNAP-2 (R. C. Bohlin et al., 2014). The two NIRSpec spectra have largely equivalent shapes, except at wavelengths shorter than about 1.4 µm: here, the spectrum measured from the 4-point nod observation drops off significantly, indicating a relative decrease in detector sensitivity toward the edges of the FOV. Meanwhile, the spectrum derived from the 20-point dither pattern is in good agreement with the CALSPEC model. It follows that the choice of one or the other SNAP-2 spectrum affects the continuum shape of the resultant reflectance spectrum, specifically the short-wavelength spectral slope.
2.5 Reflectance Spectra
2001 XR254 was situated at the nominal pointing position near the center of the FOV during our observation, and as such, we expect the wavelength-dependent sensitivity profile of the extracted spectrum to closely mirror that of the SNAP-2 spectrum measured from the 20-point dither observation. On the other hand, the location of 2016 BP81 relative to the center of the FOV in all four dither positions is more similar to the region of the detector sampled by the 4-point nod observation of SNAP-2. We therefore paired the irradiance spectra of 2001 XR254 and 2016 BP81 with the 20-point dither and 4-point nod spectra of SNAP-2, respectively, when deriving the reflectance spectra.
Lastly, we needed to account for the systematic difference between the spectra of SNAP-2 and the Sun. In Figure 2, we include the solar standard spectrum from the Planetary Spectrum Generator (PSG; G. L. Villanueva et al. 2018), which has been resampled and convolved to match the wavelength grid and spectral resolution of the NIRSpec PRISM grating. A marked discrepancy in spectral shape is seen at short wavelengths, reflecting a difference in temperature between the two stars. We corrected the reflectance spectra of our targets by multiplying the ratio between the spectra of SNAP-2 and the Sun. Figure 3 shows the final reflectance spectra of 2001 XR254 and 2016 BP81 (both components), normalized to unity at 2.5 µm.
2.6 Band Depth Calculations
To facilitate quantitative comparisons between spectra, we calculated the depths of the main absorption features of H2O and CO2 ice, which lie at 2.0, 3.0, and 4.27 µm. For each band, we stipulated narrow wavelength ranges on one or both sides of the feature that we used to establish the continuum level across the band. We also established a central wavelength region in the middle of the band that we take as the band center. Next, we carried out an error-weighted linear least-squares fit to the points within the continuum region(s) to model the local continuum. This continuum model was then divided away from the reflectance spectrum.
The band depth was computed by calculating the weighted average of all points within the band center region. The typical flux uncertainties provided by the empirical PSF fitting method in jwstspec (see Section 2.3) underestimate the true scatter of the data and cannot account for correlated noise across the wavelength axis. In order to derive more realistic band depth values and uncertainties, we calculated the ratio between the scatter in the residuals from the continuum fit and the median flux uncertainties within the continuum region(s). The flux uncertainty array of the spectrum was inflated by this ratio when calculating the band depth. The uncertainty in the band depth was set to the standard deviation of the data points within the band center region.
For the 2-µm H2O-ice absorption, the continuum regions were set to 1.80–1.90 µm and 2.20–2.30 µm, and the band center range was taken to be 2.00-2.05 µm. Only the left side of the 3-µm H2O-ice band (2.60–2.65 µm) was used to establish a flat continuum level across the broad absorption; the band depth was calculated at 2.975–3.025 µm. The CO2-ice band depth at 4.26–4.28 µm was calculated using the continuum regions 4.00–4.15 and 4.35–4.45 µm.
3 Discussion
From an initial visual inspection of the reflectance spectra, we find that the primary and secondary components of 2016 BP81 have indistinguishable spectral profiles, with a positive continuum slope throughout the 0.7–2.5 µm wavelength range, a broad 3-µm H2O ice absorption band, hints of an additional weak H2O ice feature centered at 2 µm, a deep primary CO2-ice band at 4.27 µm, and clear evidence for the 2.7-µm CO2 combination band. The strong agreement between the binary components’ spectra indicates that the two bodies have the same surface composition. This finding is in line with models of binary formation that involve gravitational collapse, in which both components accrete out of a single self-gravitating pebble cloud (e.g., A. N. Youdin & J. Goodman, 2005; D. Nesvorný et al., 2010, 2019; A. H. Parker et al., 2011). Given the identical spectral profiles of the 2016 BP81 binary components, we averaged the two together and utilize the combined spectrum in the following discussion.
In contrast, 2001 XR254 has a much steeper spectral slope below 1.0 µm, which levels off to a slightly negative-sloped continuum through 2.5 µm. At longer wavelengths, the right side of the broad 3-µm H2O-ice feature rises to significantly lower reflectance values than the left side, with a separate absorption band between 3.2 and 3.6 µm characteristic of aliphatic organics. In addition to deep CO2 features, 2001 XR254 also displays a strong CO ice absorption band centered at 4.7 µm. Minor features between 2.2 and 2.6 µm are suggestive of solid-phase CH3OH. Altogether, these distinct properties describe a surface composition very unlike that of the blue binary 2016 BP81 and is instead comparable to the CH3OH-rich compositional class of KBOs uncovered by N. Pinilla-Alonso et al. (2025). We therefore classify 2001 XR254 as a red binary cold classical object; indeed, the relatively imprecise measured visible spectral slope of %/100 nm reported in W. C. Fraser et al. (2021) straddles the threshold between the blue and red KBO subpopulations.
To examine our two targets more closely within the broader context of known KBO surface types, we reprocessed the NIRSpec observations of KBOs obtained as part of the large Cycle 1 program presented in N. Pinilla-Alonso et al. (2025), following an identical methodology to the one we applied to our cold classical targets. We excluded all of the Centaurs (inward-scattered KBOs on orbits within the giant planet region), due to the likelihood of secondary evolution across their surfaces from the enhanced temperatures and space weathering they experience; strong evidence of this trend was recently demonstrated in a dedicated analysis of their JWST spectra (J. Licandro et al., 2025). Likewise, the exceptionally H2O-ice-rich Haumea collisional family member 2003 UZ117 was not considered. The size range spanned by this spectroscope sample is 100–800 km, which overlaps at the lower end with the diameters of the cold classical targets from our program. Figure 4 shows the results of this analysis, with all spectra normalized to unity at 2.5 µm and grouped by compositional class. The combined 2016 BP81 spectrum is overplotted in each panel with the black curve, while the spectrum of 2001 XR254 is displayed in purple alongside the other CH3OH-rich objects.
From the figure, it is evident that the reflectance spectra of 2001 XR254 and 2016 BP81 generally lie within the range of spectral shapes attested across the broader CH3OH-rich and H2O-rich subclasses, respectively. The sample from N. Pinilla-Alonso et al. (2025) includes five other red cold classicals (three of them binaries), all of which belong to the CH3OH-rich compositional class, reinforcing our previous statement that 2001 XR254 is a red binary cold classical. Due to our stated focus on blue binaries in this paper, we do not consider 2001 XR254 in the following discussion. In the case of 2016 BP81, while its overall spectral profile is similar to that of the other H2O-rich KBOs, the distinctive combination of the relatively deep 4.27-µm CO2 band, the very weak 2-µm H2O-ice feature, and the absence of the 3.1-µm H2O-ice Fresnel peak appears to set it apart from the rest of its compositional class. Moreover, the reflectance level of the 2016 BP81 spectrum in the vicinity of 3.5 µm is substantially lower than the values spanned by the other spectra, suggesting that the surface of 2016 BP81 is enriched in aliphatic organics relative to the typical H2O-rich KBO.
To investigate the relationship between 2016 BP81 and other H2O-rich KBOs in a more quantitative manner, we carried out a systematic band shape analysis across the H2O-rich compositional class. Here, we included some additional H2O-rich objects from other observing programs: (1) seven Neptune Trojans from Program #2550 (PI: L. Markwardt; L. Markwardt et al. 2023), and (2) three extreme transneptunian objects (ETNOs) from Program #4665 (PI: B. Holler). All of these datasets were passed through an analogous data reduction and spectral extraction workflow. We measured the band depths of the 2- and 3-µm H2O-ice absorptions, as well as the primary CO2 band at 4.27 µm, following the methodology described in Section 2.6.
A grid plot of the two-parameter distributions is shown in Figure 5. Different symbols denote the respective dynamical subpopulations that the targets belong to, with the dynamical classifications taken from the recent work by K. Volk & C. Van Laerhoven (2024). Note that in some cases, the classifications here differ from those provided in N. Pinilla-Alonso et al. (2025).
The relationship between the 2- and 3-µm band depths is an important probe for studying variations in the relative abundance and physical properties of H2O ice across a sample of compositionally similar surfaces. Previous laboratory work (e.g., R. M. E. Mastrapa et al., 2013) and surface modeling of H2O-ice-rich solar system minor bodies (e.g., S. Protopapa et al., 2014; A. Raponi et al., 2016) have demonstrated that the two band depths vary in tandem when the overall H2O abundance changes; meanwhile, alterations to the size of H2O-ice grains manifest as changes to the relative depths between the two bands, with a higher 3- vs. 2-µm depth ratio indicative of smaller grain sizes. Looking at the top panel in Figure 5, we find that most of the H2O-rich KBOs trace out a positive correlation line, suggesting a range of relative H2O-ice abundances but otherwise similar particle size distributions.
However, 2016 BP81 is visibly offset from this covariance trend and is situated within a tight cluster alongside the Neptune Trojan 2011 SO277 and the ETNO 2016 QV89 at higher 3- vs. 2-µm depth ratios, implying smaller H2O-ice grain sizes. The spectra of these three objects are plotted together in Figure 6, from which the similarity in H2O-ice band shapes is evident. The positions of the three objects in the band-depth grid plot are linked by dashed black lines to facilitate interpretation.
When assessing the plots of the CO2 band depth as a function of the 2- and 3-µm H2O band depths in Figure 5, we once again observe a marked distinction between the blue binary 2016 BP81 and the rest of the H2O-rich KBOs. While most objects have relative CO2 band depths between 0.1 and 0.4 that are uncorrelated with their 2- or 3-µm H2O-ice absorption feature depths, both 2016 BP81 and 2011 SO277 have significantly deeper CO2-ice spectral signatures. Meanwhile, although 2016 QV89 does not stand out within the broader distribution of CO2 band depths, it nevertheless lies at the upper end of the range spanned by the other H2O-rich KBOs. From Figure 6, we can see how comparable the CO2 primary band shapes (and indeed the entire 3.5–5.1 µm spectral profiles) are among all three objects. Moreover, when examining the 3.2–3.6 µm wavelength range, we notice that 2016 QV89 exhibits a similar spectral shape to 2016 BP81, while 2011 SO277 displays an additional absorption band similar to what is seen on the CO2- and CH3OH-rich KBOs (see Figure 4). It follows that all three objects’ surfaces are enriched in aliphatic organics relative to the rest of the H2O-rich KBOs.
The discrepant spectroscopic behavior of 2016 BP81, 2011 SO277, and 2016 QV89 points toward a distinct surface subtype within the wider H2O-rich compositional class that is characterized by systematically finer H2O-ice grains, more abundant CO2 ice, and stronger signatures of aliphatic organics. We consider differences in current thermal and space weathering environment to be an unlikely explanation for this distinction, given that the three objects have drastically different present-day orbits. 2011 SO277 is particularly diagnostic in this regard, as it lies within the very narrow range of average surface temperatures and energetic particle fluences spanned by the 1:1 mean motion resonance with Neptune, while being compositionally distinct from the other six H2O-rich Neptune Trojans in our sample. Instead, we posit that the distinctive surface properties may reflect differences in the objects’ initial formation environment.
With only one blue binary in the combined spectroscopic sample, it is impossible to make general statements about the blue binary cold classical population. Observations of additional targets are necessary to determine if the distinctive properties of 2016 BP81 are a characteristic trait of the population, or if instead the blue binaries include objects with surface properties that are more typical of the broader H2O-rich KBO subclass (as is the case with the Neptune Trojans and ETNOs). The answer to that question has fundamental implications for our understanding of the origin and dynamical evolution of the blue binaries.
D. Nesvorný et al. (2022) put forward two proposals to explain the presence of blue binaries within the cold classical region (see Introduction), and we now revisit them with the observed properties of 2016 BP81 in mind. In the gentle push-out scenario, dynamical modeling has demonstrated that the survival of the most widely-separated blue binaries requires an initial source region beyond 30 au, which necessarily places the primordial blue-to-red color transition at even larger heliocentric distances. Meanwhile, the overall implantation efficiency increases with increasing initial orbital semimajor axis, which means that the current blue binary population should preferentially sample the region of the planetesimal disk nearest to the color transition. From the spectrum of 2016 BP81, we do find distinctive spectral features that could plausibly be interpreted as being transitional between those typical of the blue H2O-rich KBOs and the redder CO2- and CH3OH-rich compositional classes. These include the exceptionally strong CO2 primary band and the higher abundance of aliphatic organics, as well as the relatively weak 2-µm H2O-ice absorption (see Figure 4).
The push-out hypothesis allows for cold classical implantation from a wide range of initial heliocentric distances within the outer protoplanetary disk (e.g., D. Nesvorný et al., 2020). Therefore, one may expect future observations of blue binaries to reveal a range of surface compositions, including those that are more consistent with the majority of H2O-rich KBOs in our spectroscopic sample. Within this framework, 2011 SO277 and 2016 QV89 would also be interpreted as originating from the blue-to-red transition region prior to emplacement into their respective present-day dynamical populations. However, we recall that the simulations of the push-out scenario presented in D. Nesvorný et al. (2022) predict significant implantation of blue singletons, which is inconsistent with our current knowledge of the blue cold classical population.
Meanwhile, our spectroscopic sample presents another challenge to the push-out scenario—the apparent absence of CO2-rich objects within the cold classical population. Under the compositional gradient framework described in N. Pinilla-Alonso et al. (2025), which places the initial formation region of the CO2-rich KBOs in between that of the H2O- and CH3OH-rich subclasses, one would expect the outward implantation of the H2O-rich blue binaries from the inner transneptunian disk to be accompanied by a significant admixture of CO2-rich objects in the red cold classical population. However, all six red cold classicals with published JWST spectra (including 2001 XR254 from our observations) belong to the CH3OH-rich compositional class. In fact, out of all dynamical subpopulations within the Kuiper belt, only the cold classicals appear to lack CO2-rich objects (M. N. De Prá et al., 2025; N. Pinilla-Alonso et al., 2025). Without future refinements to the dynamical models that address the paucity of blue singletons and CO2-rich objects within the cold classical population, the push-out hypothesis appears to be untenable.
The alternative scenario described in D. Nesvorný et al. (2022) is the in situ formation of blue binaries at 42–47 au during the early stages of the protoplanetary disk, when temperatures in the disk midplane were higher and fewer volatile ices (e.g., methane and carbon monoxide) were in the solid phase (e.g., B. Bitsch et al., 2015). Objects that formed later, after the dispersion of the gaseous component of the disk and subsequent dissipative cooling of the disk midplane, would have accreted more volatile ices onto their surfaces and thereby developed the redder, CH3OH-rich surfaces typical of most cold classicals. Crucially, this proposed evolutionary pathway obviates the need to invoke dynamical implantation of the blue binaries into the cold classical region, and as such, the absence of blue singletons does not present a problem for the model. Moreover, the higher gas-to-dust ratio of the young protoplanetary disk would have facilitated the types of aerodynamic interactions that are conducive to binary formation via streaming instabilities and subsequent gravitational collapse (A. N. Youdin & J. Goodman, 2005; J. Squire & P. F. Hopkins, 2018; D. Nesvorný et al., 2021). It follows that the in situ formation scenario may be the key to explaining the near 100% binary fraction of the blue cold classicals (W. C. Fraser et al., 2017, 2021).
From a compositional perspective, the unique conditions in the outermost reaches of the young protoplanetary disk could explain the distinctive spectral features of 2016 BP81 when compared to other H2O-rich KBOs, which may have instead formed at a later time and at smaller heliocentric distances—within the warmer inner region of the transneptunian planetesimal disk. If this scenario is correct, one would expect all blue binary cold classicals to display the same characteristic compositional signatures as 2016 BP81. Consequently, the spectrally similar objects 2011 SO277 and 2016 QV89 would likely have initially formed alongside the blue binaries before being scattered from the cold classical region, perhaps due to dynamical stirring of the outer protoplanetary disk during the final phases of Neptune’s outward migration.
4 Summary
We presented JWST/NIRSpec observations of two binary cold classical KBOs. The measured reflectance spectra of the two targets differ significantly. 2001 XR254 belongs to the CH3OH-rich KBO compositional class, akin to other red cold classicals, with a steep continuum slope below 1.0 µm, strong signatures of aliphatic organics at 3.2–3.6 µm, deep features of CO2 and CO ices, and the distinctive presence of CH3OH-ice absorption bands between 2.2 and 2.6 µm. In contrast, the blue binary 2016 BP81 shows spectral features that largely coincide with the H2O-rich KBO compositional class. However, a detailed comparison of 2016 BP81 against the spectra of other H2O-rich KBOs revealed some notable differences, including a significantly stronger CO2-ice absorption feature at 4.27 µm, more prominent aliphatic organics signatures, and a higher band depth ratio between the 3- and 2-µm water-ice features. Along with the Neptune Trojan 2011 S0277 and the ETNO 2016 QV89, 2016 BP81 appears to delineate a distinct subclass within the broader H2O-rich KBO population, which may have originated in a separate formation environment from the rest of the H2O-rich KBO population.
Observations of additional blue binaries within the cold classical region are necessary to ascertain whether the unique properties of 2016 BP81 are a distinctive characteristic of this enigmatic population. Stepping back to appreciate the bigger picture, the blue binary spectrum described in this work contributes a new piece to the grand puzzle of KBO compositional diversity revealed by previous JWST observations, which have sampled almost the full range of dynamical subpopulations within the Kuiper belt. We are now in a position to leverage the compositional information provided by the body of spectroscopic data to obtain new insights into the detailed formation and evolutionary history of the Kuiper belt. Future studies that incorporate both dynamical and compositional considerations promise to greatly expand our understanding of the various physical and chemical processes that shaped the complex orbital architecture and compositional landscape of the Kuiper belt, with potentially transformative implications for our broader understanding of the outer Solar System.
This work is based on observations made with the NASA/ESA/CSA James Webb Space Telescope. The data were obtained from the Mikulski Archive for Space Telescopes at the Space Telescope Science Institute, which is operated by the Association of Universities for Research in Astronomy, Inc., under NASA contract NAS 5-03127 for JWST. These observations are associated with program #5940. The JWST data used in this paper can be found in MAST: http://dx.doi.org/10.17909/1wwx-1c65 (catalog 10.17909/1wwx-1c65).